# TSR: Trajectory‑Search Rollouts for Multi‑Turn RL of LLM Agents

### v1

@@ Abstract
Advances in large language models (LLMs) are driving a shift toward using reinforcement learning (RL) to train agents from iterative, multi-turn interactions across tasks. However, multi-turn RL remains challenging as rewards are often sparse or delayed, and environments can be stochastic. In this regime, naive trajectory sampling can hinder exploitation and induce mode collapse. We propose TSR (Trajectory-Search Rollouts), a training-time approach that repurposes test-time scaling ideas for improved per-turn rollout generation. TSR performs lightweight tree-style search to construct high-quality trajectories by selecting high-scoring actions at each turn using task-specific feedback. This improves rollout quality and stabilizes learning while leaving the underlying optimization objective unchanged, making TSR optimizer-agnostic. We instantiate TSR with best-of-$N$, beam, and shallow lookahead search, and pair it with PPO and GRPO, achieving up to 15% performance gains and more stable learning on Sokoban, FrozenLake, and WebShop tasks at a one-time increase in training compute. By moving search from inference time to the rollout stage of training, TSR provides a simple and general mechanism for stronger multi-turn agent learning, complementary to existing frameworks and rejection-sampling-style selection methods.
@@ 1 Introduction
In our experiments, we instantiate TSR with best-of-$N$, beam, and lookahead search methods, treating the corresponding search budgets (e.g., $N$, beam width, lookahead depth) as train-time compute knobs. We evaluate TSR for Qwen2.5-0.5B and Qwen2.5-3B models (Qwen et al., 2025) across three representative environments: Sokoban (Weber et al., 2018), a deterministic logic puzzle, FrozenLake (Brockman et al., 2016), a popular stochastic RL game, and WebShop (Yao et al., 2023a), a long-horizon, agentic web navigation benchmark. Under a fixed, one-time increase in train-time compute, TSR yields up to 15% absolute performance gains and significantly smoother training dynamics.
@@ 3.2 Search Strategies
We instantiate TSR with best-of-$N$, beam, and shallow lookahead search. While these are popular test-time scaling techniques traditionally applied at inference time to improve a single output, here we adapt them to training-time rollout generation for multi-turn RL.
Trajectory-Level Best-of-$N$.
For a given initial state, we independently sample a set of $N$ complete trajectories $\{\tau^{(1)},\dots,\tau^{(N)}\}$ from the base policy $\pi_{\theta}$, where each trajectory is typically scored using its cumulative reward $R(\tau)$. Rollout generation then retains only the best trajectory:
@@
Best-of-NN aims to improve training signal quality by filtering out low-reward rollouts. However, because selection occurs only after full trajectories are generated, it does not influence intermediate action choices and may still waste rollout budget on suboptimal prefixes. We therefore view trajectory-level best-of-$N$ as a naive baseline of search with depth $K$ and branching only at the root.
@@ 4.2 Training and Baselines
For our main experiments, we train Qwen2.5-0.5B and Qwen2.5-3B instruct models (Qwen et al., 2025) on Sokoban and FrozenLake, and only the larger Qwen2.5-3B model on WebShop due to the higher reasoning complexity required for this task. We use PPO for Sokoban and GRPO for FrozenLake and WebShop, as GRPO has been shown to provide more stable learning for these tasks (Wang et al., 2025). Following (Yu et al., 2025; Wang et al., 2025), we remove the KL term during optimization and apply asymmetric clipping (via clip-higher). For each batch, we sample $P=16$ task groups (prompts) and $L=16$ rollouts per group, with a maximum turn horizon of $K=5$. Our baseline is instance-level filtering with naive rollout sampling using a filtering ratio of $p=0.25$, similar to the default RAGEN implementation. When instantiating TSR with best-of-$N$, we sample $N=28$ trajectories and retain the top $L=16$. For beam search, we use a branching factor of $M=4$ and beam width $B=2$. For lookahead search, we use the same $B$ and $M$ with depth $D=2$. We selected these parameters after performing ablations on $N$, $M$, $B$, and $D$ (see App. F).
@@ 5.1 TSR Performance Comparison and Success Rates
Across all tasks and models, TSR methods consistently outperform the instance-filtering baseline, which only promotes task-level diversity but does not optimize rollouts at the per-turn level. For example, with the 3B model, TSR (Beam Search) achieves a success rate of 52.3% on Sokoban and 60.7% on FrozenLake, improvements of 8.4% and 12.0% over instance filtering. Similarly, TSR (Lookahead) and TSR (Best-of-$N$) achieve improvements ranging from 4–12%, depending on task and model size. The largest gains are observed on WebShop, where best-of-$N$, lookahead, and beam search achieve performance improvements of 7.0%, 12.3%, and 15.0%, respectively, demonstrating that TSR yields particularly strong benefits in long-horizon settings.
Among TSR variants, beam search consistently achieves the strongest performance, followed by lookahead, with best-of-$N$ providing smaller gains. These trends are also reflected in the learning dynamics, where beam search converges faster (see Fig. 3). Overall, we have the following three takeaways. First, beam search maintains multiple partial trajectories throughout rollout generation, enabling recovery from early irreversible mistakes and providing robustness to proxy-scoring noise and environment stochasticity, particularly in FrozenLake and WebShop. Second, lookahead commits to a single action based on proxy signals of future outcomes. While effective in deterministic settings like Sokoban, its gains are constrained by proxy quality and diminish as search depth increases. Third, best-of-$N$ improves training signal by selecting higher-quality completed trajectories, i.e., selection occurs only after full rollouts are generated. As a result, it acts as an oversampled variant of instance filtering, increasing the chance of “luckier” trajectories without actively steering intermediate decisions.
@@ 5.2 Exploitation, Exploration, and Training Stability
For exploitation, we analyze average training rewards in Fig. 4(a). Across all TSR variants, average rewards increase steadily over training, indicating effective learning and policy improvement. In particular, we see trends consistent with overall task performance: beam search has on average higher rewards, followed by lookahead and best-of-$N$. This indicates that TSR shifts the exploitation bias toward higher rewards, resulting in higher performance due to finding better trajectories compared to the instance filtering baseline.
@@ 6 Conclusion
In this work, we introduced Trajectory-Search Rollouts (TSR), a framework that repurposes inference-time search algorithms, such as beam search, lookahead, and best-of-$N$, to enhance the training data quality of policy gradient methods in multi-turn agentic RL. By actively exploring and pruning trajectories on a per-turn-level during the rollout phase, TSR overcomes the exploration challenges inherent in sparse- or delayed-reward environments and substantially increases final task accuracy, while effectively preventing mode collapse. Our methodology is simple, general, and optimizer-agnostic, and can thus be paired with any policy gradient method in principle. Empirically, we demonstrate that a 0.5B model trained with TSR not only significantly outperforms strong RL baselines but also surpasses larger generalist models on three representative tasks and environments. Our findings thus confirm that allocating additional compute at the trajectory rollout stage improves performance and stability in multi-turn RL training.
@@ Appendix A Related Works
Figure 5: Illustration of TSR tree-search–based rollout generation with best-of-$N$, beam search, and shallow lookahead search methods.
@@ Appendix C Trajectory Search Rollout Search Strategies
We instantiate our TSR framework with concrete search strategies inspired by test-time scaling (Snell et al., 2024), including best-of-$N$, beam search, and shallow lookahead.
@@
From a given Sokoban state, the policy samples $N$ complete rollouts independently. Trajectories that commit early to irreversible moves (e.g., pushing a box into a dead-end) terminate with low return, while trajectories that preserve maneuverability succeed. TSR retains only the highest-scoring trajectory for training.
@@
Fig. 5 provides a conceptual overview of best-of-$N$, beam search, and shallow lookahead search strategies applied in the multi-turn RL setting.
@@ Appendix E Training and Evaluation Setup
Training proceeds for up to 200 rollout–update iterations for Sokoban and FrozenLake, and 100 iterations for WebShop. At each iteration, we sample $P=16$ task instances (prompts), each generating $N=16$ rollouts, resulting in $256$ trajectories per iteration. Each trajectory is limited to a maximum of 5 interaction turns, with up to 5 actions per turn and at most 10 total actions per episode.
@@ F.2 Scaling Search Budgets
We provide additional experiments analyzing the effect of scaling search budget parameters for TSR (Beam Search) in Tables 5, 6, and 7 for Sokoban, FrozenLake, and WebShop, respectively. The results corroborate our main findings and show that the largest gains arise from increasing the beam width from $B=1$ to $B=2$. In contrast, increasing the number of action samples $M$ exhibits diminishing returns, leading to a plateauing performance trend. Overall, we find that moderate search budgets (e.g., $B=2$, $M=4$) offer a favorable trade-off between performance and computational cost. Similar trends are observed for TSR (Best-of-$N$) and TSR (Lookahead) variants, which we omit for brevity.

### v2

@@ Abstract
Advances in large language models (LLMs) are driving a shift toward using reinforcement learning (RL) to train agents from iterative, multi-turn interactions across tasks. However, multi-turn RL remains challenging as rewards are often sparse or delayed, and environments can be stochastic. In this regime, naive trajectory sampling can hinder exploitation and induce mode collapse. We propose TSR (Trajectory-Search Rollouts), a training-time approach that repurposes test-time scaling ideas for improved per-turn rollout generation. TSR performs lightweight tree-style search to construct high-quality trajectories by selecting high-scoring actions at each turn using task-specific feedback. This improves rollout quality and stabilizes learning while leaving the underlying optimization objective unchanged, making TSR optimizer-agnostic. We instantiate TSR with best-of-NN, beam, and shallow lookahead search, and pair it with PPO and GRPO, achieving up to 15% performance gains and more stable learning on Sokoban, FrozenLake, and WebShop tasks at a one-time increase in training compute. By moving search from inference time to the rollout stage of training, TSR provides a simple and general mechanism for stronger multi-turn agent learning, complementary to existing frameworks and rejection-sampling-style selection methods.
@@ 1 Introduction
In our experiments, we instantiate TSR with best-of-NN, beam, and lookahead search methods, treating the corresponding search budgets (e.g., NN, beam width, lookahead depth) as train-time compute knobs. We evaluate TSR for Qwen2.5-0.5B and Qwen2.5-3B models (Qwen et al., 2025) across three representative environments: Sokoban (Weber et al., 2018), a deterministic logic puzzle, FrozenLake (Brockman et al., 2016), a popular stochastic RL game, and WebShop (Yao et al., 2023a), a long-horizon, agentic web navigation benchmark. Under a fixed, one-time increase in train-time compute, TSR yields up to 15% absolute performance gains and significantly smoother training dynamics.
@@ 3.2 Search Strategies
We instantiate TSR with best-of-NN, beam, and shallow lookahead search. While these are popular test-time scaling techniques traditionally applied at inference time to improve a single output, here we adapt them to training-time rollout generation for multi-turn RL.
Trajectory-Level Best-of-NN.
For a given initial state, we independently sample a set of NN complete trajectories $\{\tau^{(1)},\dots,\tau^{(N)}\}$ from the base policy $\pi_{\theta}$, where each trajectory is typically scored using its cumulative reward $R(\tau)$. Rollout generation then retains only the best trajectory:
@@
Best-of-NN aims to improve training signal quality by filtering out low-reward rollouts. However, because selection occurs only after full trajectories are generated, it does not influence intermediate action choices and may still waste rollout budget on suboptimal prefixes. We therefore view trajectory-level best-of-NN as a naive baseline of search with depth $K$ and branching only at the root.
@@ 4.2 Training and Baselines
For our main experiments, we train Qwen2.5-0.5B and Qwen2.5-3B instruct models (Qwen et al., 2025) on Sokoban and FrozenLake, and only the larger Qwen2.5-3B model on WebShop due to the higher reasoning complexity required for this task. We use PPO for Sokoban and GRPO for FrozenLake and WebShop, as GRPO has been shown to provide more stable learning for these tasks (Wang et al., 2025). Following (Yu et al., 2025; Wang et al., 2025), we remove the KL term during optimization and apply asymmetric clipping (via clip-higher). For each batch, we sample $P=16$ task groups (prompts) and $L=16$ rollouts per group, with a maximum turn horizon of $K=5$. Our baseline is instance-level filtering with naive rollout sampling using a filtering ratio of $p=0.25$, similar to the default RAGEN implementation. When instantiating TSR with best-of-NN, we sample $N=28$ trajectories and retain the top $L=16$. For beam search, we use a branching factor of $M=4$ and beam width $B=2$. For lookahead search, we use the same $B$ and $M$ with depth $D=2$. We selected these parameters after performing ablations on NN, $M$, $B$, and $D$ (see App. F).
@@ 5.1 TSR Performance Comparison and Success Rates
Across all tasks and models, TSR methods consistently outperform the instance-filtering baseline, which only promotes task-level diversity but does not optimize rollouts at the per-turn level. For example, with the 3B model, TSR (Beam Search) achieves a success rate of 52.3% on Sokoban and 60.7% on FrozenLake, improvements of 8.4% and 12.0% over instance filtering. Similarly, TSR (Lookahead) and TSR (Best-of-NN) achieve improvements ranging from 4–12%, depending on task and model size. The largest gains are observed on WebShop, where best-of-NN, lookahead, and beam search achieve performance improvements of 7.0%, 12.3%, and 15.0%, respectively, demonstrating that TSR yields particularly strong benefits in long-horizon settings.
Among TSR variants, beam search consistently achieves the strongest performance, followed by lookahead, with best-of-NN providing smaller gains. These trends are also reflected in the learning dynamics, where beam search converges faster (see Fig. 3). Overall, we have the following three takeaways. First, beam search maintains multiple partial trajectories throughout rollout generation, enabling recovery from early irreversible mistakes and providing robustness to proxy-scoring noise and environment stochasticity, particularly in FrozenLake and WebShop. Second, lookahead commits to a single action based on proxy signals of future outcomes. While effective in deterministic settings like Sokoban, its gains are constrained by proxy quality and diminish as search depth increases. Third, best-of-NN improves training signal by selecting higher-quality completed trajectories, i.e., selection occurs only after full rollouts are generated. As a result, it acts as an oversampled variant of instance filtering, increasing the chance of “luckier” trajectories without actively steering intermediate decisions.
@@ 5.2 Exploitation, Exploration, and Training Stability
For exploitation, we analyze average training rewards in Fig. 4(a). Across all TSR variants, average rewards increase steadily over training, indicating effective learning and policy improvement. In particular, we see trends consistent with overall task performance: beam search has on average higher rewards, followed by lookahead and best-of-NN. This indicates that TSR shifts the exploitation bias toward higher rewards, resulting in higher performance due to finding better trajectories compared to the instance filtering baseline.
@@ 6 Conclusion
In this work, we introduced Trajectory-Search Rollouts (TSR), a framework that repurposes inference-time search algorithms, such as beam search, lookahead, and best-of-NN, to enhance the training data quality of policy gradient methods in multi-turn agentic RL. By actively exploring and pruning trajectories on a per-turn-level during the rollout phase, TSR overcomes the exploration challenges inherent in sparse- or delayed-reward environments and substantially increases final task accuracy, while effectively preventing mode collapse. Our methodology is simple, general, and optimizer-agnostic, and can thus be paired with any policy gradient method in principle. Empirically, we demonstrate that a 0.5B model trained with TSR not only significantly outperforms strong RL baselines but also surpasses larger generalist models on three representative tasks and environments. Our findings thus confirm that allocating additional compute at the trajectory rollout stage improves performance and stability in multi-turn RL training.
@@ Appendix A Related Works
Figure 5: Illustration of TSR tree-search–based rollout generation with best-of-NN, beam search, and shallow lookahead search methods.
@@ Appendix C Trajectory Search Rollout Search Strategies
We instantiate our TSR framework with concrete search strategies inspired by test-time scaling (Snell et al., 2024), including best-of-NN, beam search, and shallow lookahead.
@@
From a given Sokoban state, the policy samples NN complete rollouts independently. Trajectories that commit early to irreversible moves (e.g., pushing a box into a dead-end) terminate with low return, while trajectories that preserve maneuverability succeed. TSR retains only the highest-scoring trajectory for training.
@@
Fig. 5 provides a conceptual overview of best-of-NN, beam search, and shallow lookahead search strategies applied in the multi-turn RL setting.
@@ Appendix E Training and Evaluation Setup
Training proceeds for up to 200 rollout–update iterations for Sokoban and FrozenLake, and 100 iterations for WebShop. At each iteration, we sample $P=16$ task instances (prompts), each generating $N=16$ rollouts, resulting in 256256 trajectories per iteration. Each trajectory is limited to a maximum of 5 interaction turns, with up to 5 actions per turn and at most 10 total actions per episode.
@@ F.2 Scaling Search Budgets
We provide additional experiments analyzing the effect of scaling search budget parameters for TSR (Beam Search) in Tables 5, 6, and 7 for Sokoban, FrozenLake, and WebShop, respectively. The results corroborate our main findings and show that the largest gains arise from increasing the beam width from $B=1$ to $B=2$. In contrast, increasing the number of action samples $M$ exhibits diminishing returns, leading to a plateauing performance trend. Overall, we find that moderate search budgets (e.g., $B=2$, $M=4$) offer a favorable trade-off between performance and computational cost. Similar trends are observed for TSR (Best-of-NN) and TSR (Lookahead) variants, which we omit for brevity.